Section C: Answer any two questions ement and Question 2 ou are the audit senior of Wood LLP and you are planning the audit of Electronica Ltd (Electronica) onica specialises in manufacturing washing machines and la) Y ders o for the year ended 31 March 2017. Electr provides a two-yea with the finance director. statement extracts r warranty to its customers. Your audit manager has held a planning meeting He has provided you with the following notes of his meeting and financial Electronica has had a difficult year as household spending has fallen, hence revenue has dropped In order to address this, management has offered significantly extended credit terms to their custorners. However, demand has fallen such that there are still some machines in inventory where the selling price may be below cost. During the year, whilst calculating depreciation, the directors extended the useful lives of plant and machinery from three years to five years. This reduced the annual depreciation charge. The directors need to meet a target profit before interest and taxation of E0-5 million in order to be paid their annual bonus. In addition, to try and improve profits, Electronica changed their main material supplier to a cheaper alternative. This has resulted in more customers claiming on their warranties for repairs. To help with operating cash flow, the directors borrowed £2 million from the bank during the year. This is due for repayment at the end of 2017. Financial statement extracts for year ended 31 March DRAFT ACTUAL 2017 5m 12-5 (701 Sm 15-0 (80) Revenue Cost of sales Gross proft Operating expenses 7-0 (5-1) (50) Profit before interest and taxation 0-5 Inventory Receivables Cash Trade payables Loan 1-9 3-1 0.8 1.6 1.0 1-4 2-0 1.9 1.2 Required: Using the information above: Calculate FIVE ratios, for BOTH years, which would assist the audit senior in planning the audit; i. (10 marks) i. Using the information provided and the ratios calculated, identify and describe FIVE audit risks and explain the auditors response to each risk in planning the audit of Electronica Ltd (10 marks) Question 2 is continued on the next page
Answer #1

Part 1

Ratios

formula

2017

2016

Gross margin

Gross margin = gross profit / revenue

5.5/12.5 = 44%

7/15 = 46.7%

Operating margin

Operating margin = operating profit / revenue

0.5/12.5 = 4%

1.9/15 = 12.7%

Current ratio

Current ratio = current assets / current liabilities

5.8/1.6 = 3.6

5.3/1.2 = 4.4

Quick ratio

Quick ratio = (current assets-inventory)/current liabilities

(5.8-1.9)/1.6 = 2.4

(5.3-1.4)/1.2 = 3.3

Receivable days

Receivable days = 365* accounts receivable / sales

365*1.9/12.5 = 55 days

365*2.0/15 = 49 days

Inventory days

Inventory days = 365* inventory / cost of goods sold

365*1.9/7.0 = 99 days

365*1.4/8.0 = 64 days

Part 2

Audit Risk

Response to Risk

The results of 2017 seem to be very poor compared to 2016. There exists a risk that management may manipulate the results due to pressure on the basis of judgments take or through the use of provisions.

The audit team would constantly require to be very alert towards this risk. They have to review very carefully all judgemental decisions and analyze carefully all treatments by comparing with previous years.

There is a decrease in both gross margin and operating margin. However, the proportion of the difference between them is very high. There is a risk misclassification of costs between cost of goods sold and operating expenses.

The process of cost classification will be compared to previous years to ensure and maintain consistency.

There is a decrease in both the current and quick ratio. They both are above the standard level but still, the decrease in these ratios may result in the going concern difficulties

To ensure that the going concern basis is reasonable, the audit team should conduct the detailed going concern testing while undertaking the process of audit.

Receivables days increased from 49 days to 55 days. This shows that management has increased the credit period and thus, it presents the risk of recoverability of receivables.

To assess the valuation extended post year-end cash receipts testing should be performed. Along with it, a review of the aged receivables ledger should be conducted.

Inventory days increased from 64 days to 99 days. This increases the risk of inventory overvaluation.

The audit team needs to discuss regarding the inventory policy change with the management.